OpenStudy (anonymous):

The area of a circle is 10 pi cm. What is the circles diameter? Round to the nearest hundredth of a centimeter.

OpenStudy (anonymous):

From the problem statement: \[\pi \left(\frac{d}{2}\right)^2=10 \pi \]\[d=2 \sqrt{10} \]

OpenStudy (anonymous):

Thank you!! @robtobey

OpenStudy (anonymous):

\[N\left[2 \sqrt{10}\right]=6.32 \text{ rounded} \]

OpenStudy (anonymous):

Thank you for the medal.